# LLMs and the Human Condition

### v2

@@ 5 Representing practices
In retrospect our model is a modern take on Schank and Abelson’s scripts Schank (1989). And indeed Voice XML Voice XML was an approach to spoken language interfaces based on the notion of forms being representatives of business practices. In the latter case the community was too focused on language understanding and filling in slots in forms to notice what the forms themselves represented. The problem Schank had was that he wanted his system to represent abstractions of events so that the system could recognise two actual events as the same Schank (1972). Going to The Curry Cabin on Saturday was an instance of the going-to-a-restaurant script. Today we can represent scripts not as abstractions, but more like recordings. The enactivists talk of first person intersubjectivity in which the subject experiences interacting with the world. Rather than mapping the world into things and recognising the current situation as the same as another based on the things, the idea is that we map the situation into action and then match on the actions - actions being something the agent has direct control over. In the terminology of this paper, the environment enables a subset of the agent’s known practices, and “same” means the same enabled practice(s). Counting apples is the same as counting oranges because the process of taking apples out of a box is the same for oranges. Apples and oranges are not the same when it comes to juicing them as the physical action for jucing is different.
@@ 6 What LLMs are actually doing
In addition to “first person intersubjectivity” the enactivists also talk of “second person intersubjectivity” which is when an agent can put themselves in the shoes of a person interacting and “experience” the interaction vicariously. The mechanism for this with an embodied agent is seen as requiring some means of “body mapping” but, with a conversational agent, all that is needed is recordings.
The proposal is that the things we write down are to a very large extent descriptions of practices. Sure there are references in there to things, but LLMs don’t capture that. What an LLM is doing to generalise from N descriptions of a practice to a rewriting of the practice selected and modified by the current prompt. Prompted by “I’m hungry” an LLM will “select” the practice of mother feeding a child - because this is a standard opening for a child to a mother - and fill it in. If the prompt includes the statement that the child likes cereal, then “want some cereal” would be a very plausible magic response.

### v3

@@ 5 Representing practices
In retrospect our model is a modern take on Schank and Abelson’s scripts Schank (1989). And indeed Voice XML Voice XML was an approach to spoken language interfaces based on the notion of forms being representatives of business practices. In the latter case the community was too focused on language understanding and filling in slots in forms to notice what the forms themselves represented. The problem Schank had was that he wanted his system to represent abstractions of events so that the system could recognise two actual events as the same Schank (1972). Going to The Curry Cabin on Saturday was an instance of the going-to-a-restaurant script. Today we can represent scripts not as abstractions, but more like recordings. The enactivists talk of first person intersubjectivity in which the subject experiences interacting with the world. Rather than mapping the world into things and recognising the current situation as the same as another based on the things, the idea is that we map the situation into action and then match on the actions - actions being something the agent has direct control over. In the terminology of this paper, the environment enables, or affords, a subset of the agent’s known practices, and “same” means the same enabled practice(s). Counting apples is the same as counting oranges because the process of taking apples out of a box is the same for oranges. Apples and oranges are not the same when it comes to juicing them as the physical action for juicing is different.
@@ 6 What LLMs are actually doing
In addition to “first person intersubjectivity” the enactivists also talk of “second person intersubjectivity” which is when an agent can put themselves in the shoes of a person interacting and “experience” the interaction vicariously. Observing someone cook gnocchi and then at some time in the future cooking it oneself, the conjectured mechanism does not remember pots, stoves and water, nor reason about why he put the lid on the pot, but simply “silently” copies the actions of the cook while the cook is cooking. These vicarious actions are then replayed when it comes time to cook gnocchi. The key is that abstraction of sense data to things such as pot lids such as it is , is done by “mirror neurons” or something similar. What is remembered is first person actions mirrored from the second person experience.
Applied to conversation the mapping from observed action of a second party to one own speech-acts is very straight-forward: one simply says the words that one heard someone else say in that situation. For disembodied conversation such as phone calls or text on the internet, all that is needed is recordings or text. This is of course what LLMs are trained on.
The proposal is that the things we write down are to a very large extent descriptions of practices. Sure there are references in there to things, but LLMs don’t capture that. A Generative Predictive Transformer – or any mechanism for generative prediction – is not “understanding” that a cat has four legs, sharp teeth and fur, what it is learning is that we pat cats, we attend when they catch mice, and that they jump on window sills and knock things off. The things we notice about cats are all integrated into our own practices, and what we talk about are those practices.
What an LLM is doing to generalise from N descriptions of a practice to a rewriting of the practice selected and modified by the current prompt. Prompted by “I’m hungry” an LLM will “select” the practice of mother feeding a child - because this is a standard opening for a child to a mother - and fill it in. If the prompt includes the statement that the child likes cereal, then “want some cereal” would be a very plausible magic response.